OpenStudy (anonymous):

the integral of 4/(x^2 +4) from -10 to 10

OpenStudy (anonymous):

Have you learned Trig Sub yet?

OpenStudy (anonymous):

I don't know...what is it? I am using integration and limits in this section

OpenStudy (anonymous):

you know, limits of integration and such

OpenStudy (anonymous):

What are some of the method of integration you have learned?

OpenStudy (anonymous):

rule, algerbra, u-sub, integration by parts, trigonometricl integrals, improper integrals

OpenStudy (anonymous):

We may rewrite you integrand as following :\[4\over 4({1 \over 4}x^2+1)\]

OpenStudy (anonymous):

Also I have learned indeterminate forms

OpenStudy (anonymous):

canceling out 4s we get this \[1 \over {1 \over4}x^2+1\]

OpenStudy (anonymous):

There is a general form: \[\int\limits \frac{du}{u^2+a^2}=\frac{1}{a}\arctan(\frac{u}{a})+C\] So the integral is just: \[4*\frac{1}{2}\arctan(\frac{x}{2})+C=2\arctan(\frac{x}{2})+C\]

OpenStudy (anonymous):

Yeas, I have used that one, but the answer I received from the problem, (1/2)arctan(5) is incorrect

OpenStudy (anonymous):

Then you have your bounds. -10 to 10 You have: \[2(\arctan(-5)-\arctan(5))\]

OpenStudy (anonymous):

Then you can get an approximation if you want or leave it as exact.

OpenStudy (anonymous):

So I cannot use the bounds 0 to 10, and then multiply by two?

OpenStudy (anonymous):

Not necessarily. That only works if it is an even function.

OpenStudy (anonymous):

and arctan is odd I believe.

OpenStudy (anonymous):

It is. :P So you can't multiply it by 2.

OpenStudy (anonymous):

oh....

OpenStudy (anonymous):

So you have 2(-arctan(5)-arctan(5))=-4arctan(5). Since arctan is odd arctan(-x)=arctan(x)

OpenStudy (anonymous):

but 2(arctan(-5)-arctan(5)) doesn't work either. I put in -4arctan(5) also before coming to this site and that doesn't work either.

OpenStudy (anonymous):

neither does 4arctan(5)

OpenStudy (anonymous):

Before I came here I also put that answer in with no such luck. Here is a clip of what I am doing.

OpenStudy (anonymous):

You were right, in this case you can do 0 to 10 and multiply by 2. Because 1/x^2+4 is even. (I jumped ahead of myself) Either way, you should still get the same answer. Do they want an approximation?

OpenStudy (anonymous):

It doesn't say, but when I used wolfram's approximation, it's labeled incorrect

OpenStudy (anonymous):

Then I really don't know what to tell you man :/ We did the problem correct.

OpenStudy (anonymous):

That's what I think too. Guess it is a mistake in the system. I will ask my teacher about it. Thanks for helping!

OpenStudy (anonymous):

No problem. Sorry for the misthought on the even/odd. You don't look at the integral you look at the original function xP